OpenStudy (anonymous):

Simplify 16 divided by 2 + 6 – 4 x 3

OpenStudy (anonymous):

\(\large 16 \div 2 + 6 - 4 * 3\) or \(\huge \frac{16}{2+6-4*3}\)

OpenStudy (anonymous):

wouldn't be -4

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

If it's the first thing I typed, yes it comes out to 2. Order of operations says you should do the division and multiplication before the addition and subtraction.

OpenStudy (anonymous):

\(\Large 16 \div 2 +6 -4*3 = 8+6-12 = 14-12 = 2\)